# Predicting post-release defects with knowledge units (KUs) of programming languages: an empirical study

### v1

@@ Abstract
Traditional code metrics (product and process metrics) have been widely used in defect prediction. However, these metrics have an inherent limitation: they do not reveal system traits that are tied to certain building blocks of a given programming language. We conjecture that taking these building blocks of a programming language into account can lead to further insights about a software system and improve defect prediction. Towards filling this gap, this paper reports an empirical study on the usage of knowledge units (KUs) of the Java programming language to predict post-release defects. A KU is a cohesive set of key capabilities that are offered by one or more building blocks of a given programming language. We operationalize our KUs via certification of the Java programming language. This study aims to understand whether we can obtain richer results in defect prediction (e.g., more accurate prediction of defect-prone code) when using KUs in combination with traditional code metrics. We select the defect dataset that was curated by (Yatish et al., 2019) which includes defect dataset for 28 releases of 8 Java software systems. We analyze the source code of these releases and extract traditional code metrics and the incidence of KUs. We find empirical evidence that KUs are different and complementary to traditional metrics, thus indeed offering a new lens through which software systems can be analyzed. Such a result motivates us to study the effectiveness of KUs in predicting defect-prone code. We build a defect prediction model called KUCLS, which leverages the KU-based features. Our KUCLS achieves a median AUC of 0.82 and significantly outperforms the CC_PROD (model built with product metrics). The normalized AUC improvement of the KUCLS over CC_PROD ranges from 5.1% to 28.9% across the studied releases. Although the KUCLS does not outperform the CC model (model built with product and process metrics), it maintains strong AUC values across the studied releases. Combining KUs with traditional code metrics leads to even higher-performing models (KUCLS+CC), significantly outperforming both the CC (an improvement ranges from 4.9% to 33.3%) and KUCLS models (an improvement ranges from 5.6% to 59.9%). Finally, we develop a cost-effective model (KUCLS_CC_COST_EFF) that significantly outperforms the CC. These encouraging results can be helpful to researchers who wish to further study the aspect of feature engineering and building models for defect prediction. Future work in this area thus (i) consider the KUCLS as a baseline model and (ii) include KU-based features in the design of models that predict post-release defects.
@@
Software engineering , Software metrics , code metrics , Java , certification exam , defect prediction , machine learning
[…]
[1] organization=Software Analysis and Intelligence Lab (SAIL), School of Computing, Queen’s University, city=Kingston, Ontario, country=Canada
[…]
[2] organization=York University, city=Toronto, Ontario, country=Canada
@@ 1 Introduction
Defect prediction has been a long-standing area of research in the field of software engineering, aimed at identifying potential defect-prone code in software systems (Koru et al., 2008; Nagappan et al., 2006; Moser et al., 2008; Rahman and Devanbu, 2013; Majumder et al., 2022). While the defect prediction field has seen substantial research in developing new machine learning approaches (e.g., bootstrapped by the recent advances and popularization of deep learning techniques (Hoang et al., 2019; Li et al., 2017; Chen et al., 2019; Zain et al., 2023)), the exploration of novel types of data for feature engineering remains less explored. The continual improvement in the defect prediction field demands innovation not just in modeling techniques but also in the data used for feature engineering. The more indicators researchers have, the more ammunition they have to understand defects and conjecture their real causes.
@@
Traditional code metrics (e.g., product and process metrics) have been extensively used for defect prediction over several decades. These metrics characterize the size (e.g., LOC), complexity (e.g., cyclomatic complexity), and structure (e.g., the CK suite (Chidamber and Kemerer, 1994)) of object-oriented software systems. These metrics are programming-language agnostic, in the sense that their definitions are applicable to and computable for any object-oriented system (irrespective of the used programming language). As an inherent limitation, traditional code metrics do not reveal system traits that are tied to certain building blocks of a given programming language. As an illustrative example, let us consider a system written in Java and that makes use of the Java Concurrency API (i.e., a few classes from the java.util.concurrent package). Detecting the use of such an API reveals that concurrent code is a trait of that system. Such a trait is relevant: specific defects might arise because of it, not every developer is proficient in concurrent programming, and debugging a concurrent program is often more difficult than debugging a single-threaded one. Yet, traditional code metrics would not capture this concurrency trait.
@@
Towards filling this gap, this paper reports an empirical study on the usage of knowledge units (KUs) of the Java programming language to predict post-release defects. As defined in our prior study (Ahasanuzzaman et al., 2024) a KU is a cohesive set of key capabilities that are offered by one or more building blocks of a given programming language (see Figure 1). The word “capabilities” refer to the main things that a programmer can do with a certain building block of a programming language. For instance, using the Concurrency API building block, a developer can create worker threads to execute tasks concurrently. Therefore, it is reasonable to assume that Java has a Concurrency KU, which includes a cohesive set of key concurrent program capabilities offered by the Concurrency API (building block). We operationalize our KUs using certification exams for the Java programming language.
In this paper, we conduct an exploratory study on the KUs of the Java programming language for predicting post-release defects. Our goal is to understand whether we can obtain richer results in defect prediction (e.g., more accurate prediction of defect-prone code) when using KUs in combination with traditional code metrics. We proceed as follows. First, we select the defect dataset that was curated by (Yatish et al., 2019). This dataset contains post-release defect data for 28 releases that span a total of 8 Java software systems. Then, we analyze the source code of each of the studied releases and collect the 28 KU-based features from each file of the defect dataset using our custom built KU detector. We also extract 59 traditional code metrics (54 product metrics and 5 process metrics) that have been used for defect prediction in prior studies. We start with a preliminary study to compare KU-based features with traditional code metrics. We apply the K-means algorithm to cluster the files of the studied releases using KUs. Subsequently, we recluster the systems using traditional code metrics and compare the two clusterings. We observe that KUs identify two times more cluster than traditional code metrics and the identified clusters of KUs do not overlap with the clusters of code metrics. Therefore, KUs are different from code metrics and can offer a new lens to analyze the source code of software systems. This study motivates us to study the effectiveness of KUs in post-release defect prediction. More specifically, we address the following six research questions (RQs):
@@
RQ1: How effective are KU-based features for identifying defect prone code? First, we use the KU-based features to build a defect prediction model for each of the studied releases. We refer to this model as KUCLS. To compare the performance of the KUCLS, we build two baseline models using traditional code metrics: (i) CC_PROD model which is built with product metrics and (ii) CC model which is built with all the studied traditional code metrics (both process and product metrics). To compare the performance of KUCLS and any of the baseline model, we perform a statistical test using a two-sided Wilcoxon signed-rank test. Additionally, we employ Cliff’s delta (Long et al., 2003) as an effect size measure to assess the practical significance of the differences between the distributions. Our results indicate that:
@@
KUCLS achieves a median AUC of 0.82. A defect prediction model with an AUC of 0.70 or higher is typically considered suitable for practical use (Jiarpakdee et al., 2020). The KUCLS outperforms the CC_PROD (median AUC of 0.80) in the defect prediction task. The normalized AUC improvement of the KUCLS over the CC_PROD ranges from 5.1% to 28.9% across the studied releases. Although the KUCLS does not outperform the CC model (median AUC of 0.84), it maintains strong AUC values across the studied releases.
@@
RQ2: What are the important KU-based features of the KUCLS for predicting post-release defects? To determine the feature dimensions that significantly impact the KUCLS’s performance, we utilize SHAP (SHapley Additive exPlanation). SHAP is a robust, flexible, and widely used method for model interpretation (Molnar, 2020; Rajbahadur et al., 2022). We first compute the absolute SHAP values for each KU-based feature across all records in the dataset. Next, for each feature, we sum the SHAP values of all records in the dataset. This process is repeated for 100 bootstrapped datasets per release, resulting in a distribution of SHAP value for each feature. We then apply the Scott-Knott ESD (SK-ESD) technique (Ghotra et al., 2015; Mittas and Angelis, 2013; Tantithamthavorn et al., 2017) to rank features based on these distributions and obtain a feature ranking for each release. We take the ranking of a feature for all releases and produce a feature rank distribution for the feature. To summarize the results across all releases, we apply the Scott-Knott ESD technique again on the feature rank distributions, similarly to Ghotra et al. (2015). Our results indicate that:
@@
Method Encapsulation, Inheritance, and Exception are identified as the top three most influential KU-based features of the KUCLS for predicting post-release defects consistently across all studied releases.
RQ3: Can KU-based features of the KUCLS provide insights into the incidence of post-release defects? To understand detailed insights into the factors influencing predictions for individual code segments, we perform an instance specific interpretation (a.k.a., local interpretations) using SHAP. We manually investigate a few buggy files, their corresponding issues in an ad-hoc manner and the SHAP values of the features of KUCLS and CC. Our results indicate that:
KU-based features offer insights into the characteristics of defect-prone files. Such insights are complementary to those offered by traditional code metrics.
@@
RQ4: Can KUCLS be made more accurate by combining KU-based features with traditional code metrics? We combine the KU-baed features of the KUCLS and the features of the CC to determine whether we can obtain a higher-performing model. More specifically, we use all KU-based features of the KUCLS and all traditional code metrics features of the CC to build a combined model, which we call KUCLS+CC. We then apply the SK-ESD to compare the performance of KUCLS+CC with that of KUCLS and CC. To investigate the feature dimension of KUCLS+CC that most significantly influence predictions, we employ the same SHAP method from RQ2. Our results indicate that:
@@
KUCLS+CC achieves a median AUC of 0.87 and outperforms both the KUCLS and CC across all the studied releases. The normalized AUC improvement of the KUCLS+CC over the CC ranges from 4.9% to 33.3% and the improvement over the KUCLS ranges from 5.6% to 59.9% across the studied releases. Although DDEV (a code metric feature) emerges as the most influential feature overall, Inheritance and Method Encapsulation KU-based features are the third ranked features indicating significant contribution of KU-based feature in predicting post-release defects.
RQ5: How different classification algorithms with hyper-parameter tuning impact the combined model for predicting post-release defects? We build the KUCLS+CC using a random forest classifier with default hyper-parameter values (scikit-learn python implementation). Since each classifier has its own strengths and weaknesses, the optimal performance of a prediction model can be achieved by choosing an appropriate classifier and tuning its hyper-parameters. To explore how different classifiers with hyper-parameter tuned parameters affect the KUCLS+CC model, we evaluate a mix of traditional and advanced classifiers: Naive Bayes (NB), k-nearest neighbor (KNN), Decision Tree (DT), Random Forest (RF), XGBoost (XGB), and LightGBM (LGBM). We build the KUCLS+CC models using each of these selected classifiers. To determine the most effective model hyper-parameter configuration, we utilize Scikit-learn’s GridSearchCV [^2] https://scikit-learn.org/stable/modules/generated/sklearn.model_selection.GridSearchCV.html. We compare these models and rank them using SK-ESD technique. Our results indicate that:
Among the studied classifiers, the KUCLS+CC that is built using a random forest classifier with default hyper-parameter values ranks as the highest performing model. Hence, we cannot further improve the performance of the KUCLS+CC from RQ4 by leveraging other classifiers and tuned hyper-parameters.
@@
RQ6: How does the performance of a cost-effective combined model compare to that of the model built with traditional code metrics? A cost-effective model that uses fewer features yet maintains decent performance is thus valuable, as it reduces operational burdens associated with extensive feature engineering. Towards this goal, we build a cost-effective combined model, which we call KUCLS_CC_COST_EFF. This model is built using the top five features of the KUCLS and the top five code metric features of the CC. This results in a combined model with only 10 features, compared to the 59 features in CC and the 28 features in KULTC. Our results indicate that:
@@
The KUCLS_CC_COST_EFF achieves a median AUC of 0.86 and outperforms the CC. The normalized AUC improvement of the KUCLS_CC_COST_EFF over the CC ranges from 5.5% to 33.3%. Therefore, we improve over the model built with traditional code metrics while still keeping feature engineering cost at a reasonable level.
@@
Paper organization. Section 2 defines KUs and presents our approach for detecting them. Section 3 describes our data collection approach. Section 5 presents the motivation, approach, and findings that are associated with our research questions. Section 6 discusses the balance between performance improvement vs model complexity and directions to improvements of models. Section 7 describes the threats to the validity of our findings. Section 8 discusses related work. Finally, Section 9 concludes the paper.
@@ 2.2 Eliciting KUs
We reuse the KUs that we elicited in our prior work (Ahasanuzzaman et al., 2024). To elicit those KUs, we relied on the Oracle certification exams for the Java programming language. Oracle offers certification exams for different Java editions, such as Java Standard Edition (Java SE) and Java Enterprise Edition (Java EE). Specifically, for Java SE, there are two levels of certification exams:(i) Java SE 8 Programmer I Certification exam (Oracle, 2022a), and (ii) Java SE 8 Programmer II Certification exam(Oracle, 2022b). For Java EE, only one version of the exam is offered by Oracle: Oracle Certified Professional, Java EE Application Developer Certification exam (Oracle, 2022c). We elicited KUs directly from the topics outlined in these exams. In most cases, an exam topic was interpreted as a KU (e.g., the String Processing topic of Java SE 8 Programmer II Certification exam was interpreted as “Generics and Collections”) and its subtopics (e.g., create a generic class, use TreeSet, TreeMap and ArrayList and use java.util.Comparator) were interpreted as the key capabilities of that KU. At the end of this process, we elicited 28 KUs (Table 1). We refer to our prior work (Ahasanuzzaman et al., 2024) for the detailed description of this mapping process.
@@
[K2] Operator and Decision The usage of different Java operators (e.g., assignment, logical, and bit-wise operators) and conditional statements (e.g, if, if-else, and switch statements).
@@
[K5] Method and Encapsulation The creation of methods with parameters, the use of overloaded methods and constructors, the usage of constructor chaining, the creation of methods with variable length arguments, and the usage of different access modifiers. This KU also describes encapsulation mechanisms, such as creating a set and get method for controlling data access, generating immutable classes, and updating object type parameters of a method.
@@
[K8] Generics and Collection The creation and usage of generic classes, usage of different types of interfaces (e.g., List Interface, Deque Interface, Map Interface, and Set Interface), and comparison of objects using interfaces (e.g., java.util.Comparator, and java.lang.Comparable).

### v2

# ✉ Zhen Ming (Jack) Jiang

@@ Abstract
Defect prediction plays a crucial role in software engineering, enabling developers to identify defect-prone code and improve software quality. While extensive research has focused on refining machine learning models for defect prediction, the exploration of new data sources for feature engineering remains limited. Defect prediction models primarily rely on traditional metrics such as product, process, and code ownership metrics, which, while effective, do not capture language-specific traits that may influence defect proneness. To address this gap, we introduce Knowledge Units (KUs) of programming languages as a novel feature set for analyzing software systems and defect prediction. A KU is a cohesive set of key capabilities that are offered by one or more building blocks of a given programming language. We conduct an empirical study leveraging 28 KUs that are derived from Java certification exams and compare their effectiveness against traditional metrics in predicting post-release defects across 28 releases of 8 well-maintained Java software systems. Our results show that KUs provide significant predictive power, achieving a median AUC of 0.82, outperforming individual group of traditional metric-based models (e.g., process, product and ownership metrics). Among KU features, Method & Encapsulation, Inheritance, and Exception Handling emerge as the most influential predictors. Furthermore, combining KUs with traditional metrics enhances prediction performance, yielding a median AUC of 0.89. We also introduce a cost-effective model using only 10 features (5 KUs and 5 traditional metrics), which maintains strong predictive performance while reducing feature engineering costs. Our findings demonstrate the value of KUs in predicting post-release defects, offering a complementary perspective to traditional metrics. This study can be helpful to researchers who wish to analyze software systems from a perspective that is complementary to that of traditional metrics.
@@
Defect prediction , Software metrics, code metrics, Java, certification exam, software engineering, and machine learning
@@ 1 Introduction
Defect prediction has been a long-standing area of research in the field of software engineering, aimed at identifying potential defect-prone code in software systems (Koru et al., 2008b; Nagappan et al., 2006b; Moser et al., 2008; Rahman and Devanbu, 2013; Majumder et al., 2022). While the defect prediction field has seen substantial research in developing new machine learning approaches (e.g., bootstrapped by the recent advances and popularization of deep learning techniques (Hoang et al., 2019; Li et al., 2017; Chen et al., 2019; Zain et al., 2023), the exploration of novel types of data for feature engineering remains less explored. The continual improvement in the defect prediction field demands innovation not just in modeling techniques but also in the data used for feature engineering. The more indicators researchers have, the better equipped they are to understand defects and hypothesize their underlying causes.
@@
Product, process and code ownership metrics have been extensively used for defect prediction over several decades. We collectively refer to them as traditional metrics. These metrics characterize the size (e.g., LOC), complexity (e.g., cyclomatic complexity), and structure (e.g., the CK suite (Chidamber and Kemerer, 1994)) of object-oriented software systems. Most importantly, these metrics are programming-language agnostic, in the sense that their definitions are applicable to and computable for any object-oriented system (irrespective of the used programming language). As an inherent limitation, traditional metrics do not reveal system traits that are tied to certain building blocks of a given programming language. As an illustrative example, let us consider a system written in Java and that makes use of the Java Concurrency API (i.e., a few classes from the java.util.concurrent package). Detecting the use of such an API reveals that concurrent code is a trait of that system. Such a trait is relevant: specific defects might arise because of it, not every developer is proficient in concurrent programming, and debugging a concurrent program is often more difficult than debugging a single-threaded one. Yet, traditional metrics would not capture this concurrency trait.
@@
Towards filling this gap, this paper reports an empirical study on the usage of knowledge units (KUs) of the Java programming language for predicting post-release defects. As defined in our prior study (Ahasanuzzaman et al., 2024), a KU is a cohesive set of key capabilities that are offered by one or more building blocks of a given programming language (see Figure 1). The word “capabilities” refer to the main things that a programmer can do with a certain building block of a programming language. For instance, using the Concurrency API building block, a developer can create worker threads to execute tasks concurrently. Therefore, it is reasonable to assume that Java has a Concurrency KU, which includes a cohesive set of key concurrent program capabilities offered by the Concurrency API (building block). Our hypothesis is that by capturing system traits such as concurrency through KUs, we will be able to design more accurate post-release defect prediction models compared to those relying exclusively on traditional metrics.
We note that KUs are also able to capture more nuanced code characteristics (e.g., compared to concurrent programming) that may still influence defect-proneness. For instance, a method is a fundamental building block in Java. Developers can create methods with parameters, overloaded methods, methods with variable-length arguments, and define access modifiers for methods. Not every object-oriented language supports these capabilities. Therefore, it is also reasonable to assume that Java has a Method & Encapsulation KU, which captures the key capabilities associated with methods and constructors. KUs capture deeper concepts than merely counting keywords. For instance, detecting the Method & Encapsulation KU involves tracking the usage of variable length arguments in methods and the application of overloaded constructors and constructor chaining. To achieve this, we use the Eclipse JDT framework to parse the code’s Abstract Syntax Tree (AST), allowing us to gather nuanced and richer information that goes much beyond simple keyword matching. By incorporating such granular details, KUs offer a more comprehensive perspective on software structure that extends beyond traditional metrics.
As described in our prior work (Ahasanuzzaman et al., 2024), we elicit KUs by manually analyzing three certification exams for the Java programming language (Section 2.2): Java SE 8 Programmer I Exam, Java SE 8 Programmer II Exam, and Java EE Developer Exam. We elicit a total of 28 KUs, which cover a wide range of capabilities (A). To count the occurrences of each KU in a Java source code file, we built a custom KU detector on top of the Eclipse JDT framework (Section 2.3). We refer to the set of metrics corresponding to the counts of each KU as KU metrics.
Our data collection is designed as follows. First, we select the defect dataset that was curated by Yatish et al. (2019). This dataset contains post-release defect data for 28 releases that originate from 8 Java systems. For each source code file in a release, we compute our 28 KU metrics. Additionally, we reuse all 65 traditional metrics provided in the defect dataset, including 54 product metrics, 5 process metrics, and 6 code ownership metrics for each source code file. Finally, we build defect prediction models using these two sets of metrics to address the following research questions (RQs):
@@
RQ1: How effective are KU features for predicting post-release defects? First, we use the KU metrics to build a defect prediction model for each of the studied releases. We refer to this model as KUM. To judge the performance of KUM, we build four baseline models using traditional metrics: (i) PROD (which is built with product metrics), (ii) PROC (which is built with process metrics), (iii) OWN (which is built with code ownership metrics) and (iv) TM (which is built using all product, process and code ownership metrics). To compare the performance of KUM and any of the baseline models, we perform a statistical test using a two-sided Wilcoxon signed-rank test. Additionally, we employ Cliff’s delta (Long et al., 2003) as an effect size measure to assess the practical significance of the differences between the distributions. We also rank the studied models using the Scott-Knott ESD (SK-ESD) technique (Ghotra et al., 2015; Mittas and Angelis, 2013; Tantithamthavorn et al., 2017). Our results indicate that:
@@
KUM achieves a median AUC of 0.82 (a defect prediction model with an AUC of 0.70 or higher is typically considered suitable for practical use). KUM outperforms three baseline models: PROD, PROC and OWN. The normalized AUC improvement of KUM over PROD ranges from 3.6% to 22.2% (median 11.4%), improvement over PROC ranges from 6.7% to 68.9% (median 42.4%), and improvement over OWN ranges from 3.3% to 71.9% (median (33.3%). That is, KUM not only outperforms a model built on the same raw data (PROD), but also those built on process (PROC) and ownership (OWN) data. TM is the top-performing model, achieving a median AUC of 0.85
@@
RQ2: What are the most important KU features for predicting post-release defects? To determine the most important features of KUM, we utilize SHAP (SHapley Additive exPlanation). SHAP is a robust, flexible, and widely used method for model interpretation (Molnar, 2020; Rajbahadur et al., 2022). Next, we apply the SK-ESD technique on the distribution of SHAP values to rank features. Our results indicate that:
@@
Method & Encapsulation, Inheritance, and Exception are identified as the top three most important KU features of KUM for predicting post-release defects.
@@
RQ3: Can KUM be made more accurate by combining its features with those from TM? Based on the encouraging results from RQ1, we join all the features of KUM and TM to build a combined model, which we call KUM+TM. We compare the performance of KUM+TM with that of KUM and TM using the same approach from RQ1. To investigate the importance of feature dimensions in KUM+TM, we employ the same SHAP approach from RQ2. Our results indicate that:
@@
KUM+TM achieves a median AUC of 0.89 and outperforms both KUM and TM across all of the studied releases. The normalized AUC improvement of KUM+TM over TM ranges from 11.1% to 44.4% (median 22.5%) and the improvement over KUM ranges from 10.5% to 72.2% (median 31.5%) across the studied releases. The most important feature is the number of active developers (ADEV, a process feature). The top ranked KU feature is Method & Encapsulation at rank 4. Inheritance and Exception Handling KU features are also among the top ten most important features.
@@
RQ4: Can a cost-effective model be created by combining specific features from KUM and TM? A cost-effective model that uses fewer features and yet maintains decent performance is valuable, as it reduces operational costs associated with extensive feature engineering. Towards this goal, we build a cost-effective combined model, which we call COST_EFF. This model is built using the top five features of KUM and the top five traditional metric features of TM. This results in a combined model with only 10 features, compared to the 65 features in TM and the 28 features in KULTC. Our results indicate that:
@@
The COST_EFF achieves a median AUC of 0.87 and outperforms TM. The normalized AUC improvement of COST_EFF over TM ranges from 3.7% to 31.3% (median 11.1%). Therefore, we improve over the model built with all traditional metrics while still keeping feature engineering cost at a much more reasonable level.
The main contribution of this paper is to evaluate the effectiveness of KUs of the Java programming language for predicting post-release defects. We believe that our findings can be helpful and encouraging to researchers who wish to analyze software systems from a perspective that is complementary to traditional metrics. Our encouraging results indicate that further refining the concept of KUs and exploring alternative elicitation techniques can potentially lead to even higher-performing defect prediction models. To bootstrap future work in this area, we provide online supplementary material with the data analyzed as part of this study [^1] http://bit.ly/3GoSmHL.
@@
Paper organization. Section 2 defines KUs and presents our approach for detecting them. Section 3 describes our data collection approach. Section 4 presents the motivation, approach, and findings that are associated with our research questions. Section 5 discusses the trade-offs between traditional metrics and KUs for defect prediction and future directions to expanding KUs. Section 6 describes the threats to the validity of our findings. Section 7 discusses related work. Finally, Section 8 concludes the paper.
@@ 2.2 Eliciting KUs
We reuse the KUs that we elicited in our prior work (Ahasanuzzaman et al., 2024). To elicit those KUs, we relied on the Oracle certification exams for the Java programming language. Oracle offers certification exams for different Java editions, such as Java Standard Edition (Java SE) and Java Enterprise Edition (Java EE). Specifically, for Java SE, there are two levels of certification exams:(i) Java SE 8 Programmer I Certification exam (Oracle, 2022c), and (ii) Java SE 8 Programmer II Certification exam(Oracle, 2022a). For Java EE, only one version of the exam is offered by Oracle: Oracle Certified Professional, Java EE Application Developer Certification exam (Oracle, 2022b). We elicited KUs directly from the topics outlined in these exams. In most cases, an exam topic was interpreted as a KU (e.g., the Generics and Collections topic of Java SE 8 Programmer II Certification exam was interpreted as “Generics and Collections”) and its subtopics (e.g., create a generic class, use TreeSet, TreeMap and ArrayList and use java.util.Comparator) were interpreted as the key capabilities of that KU. At the end of this process, we elicited 28 KUs (Table 1). We refer to our prior work (Ahasanuzzaman et al., 2024) for the detailed description of this mapping process.
@@
[K2] Operator & Decision The usage of different Java operators (e.g., assignment, logical, and bit-wise operators) and conditional statements (e.g, if, if-else, and switch statements).
@@
[K5] Method & Encapsulation The creation of methods with parameters, the use of overloaded methods and constructors, the usage of constructor chaining, the creation of methods with variable length arguments, and the usage of different access modifiers. This KU also describes encapsulation mechanisms, such as creating a set and get method for controlling data access, generating immutable classes, and updating object type parameters of a method.
@@
[K8] Generics & Collection The creation and usage of generic classes, usage of different types of interfaces (e.g., List Interface, Deque Interface, Map Interface, and Set Interface), and comparison of objects using interfaces (e.g., java.util.Comparator, and java.lang.Comparable).